# Network Modelling of Asynchronous Change-Points in Multivariate Time Series

### v1

@@ 4.1 Simulation Study
Table 2: Average log Bayes Factors, $log\text{BF}_{M}$, and 5% quantiles under each scenario for Gaussian mean likelihood. Highest average Bayes Factor for each scenario is marked in bold.
Table 2 displays the averaged log Bayes factors, relative to NetCP for each model under the normal means likelihood scenario. In S1 and S2, which involve lead-lag relationships in change-points, NetCP clearly outperforms all other methods, with substantial negative log Bayes factors observed for competing models. In S3, where change-points occur simultaneously across series, NG-CCP achieves the highest score relative to NetCP, though the interval overlaps zero, suggesting comparable performance. In S4, G-CCP performs best, though differences between methods are again modest. For S5, which contains independent change-points across series, G-CCP yields the highest average Bayes factor, but Yao performs similarly, as indicated by their overlapping intervals. Overall, these results suggest that NetCP excels in structured, temporally misaligned settings, while G-CCP and Yao are more competitive in scenarios with simultaneous or independent change-points.
@@
Table 3: Average log Bayes Factors, $log\text{BF}_{M}$, and 5% quantiles under each scenario for AR process likelihood. Highest average Bayes Factor for each scenario is marked in bold.
@@ 4.2 Seismology Data
For our first real example, we apply our model to the detection of microearthquakes using data from a network of seismic sensors. Specifically, we use recordings from the High Resolution Seismic Network (HSRN), maintained by the UC Berkeley Seismology Laboratory (NCEDC,, 2014). This network comprises geophone sensors positioned around the town of Parkfield, California, and is designed to monitor seismic activity along the San Andreas fault. Similar data was previously analysed by Xie et al., (2019) who took a frequentist approach, focussing on estimating the relative delay between change-points in different series. For our analysis, we consider data from seven sensors over a 100 second period beginning at 22:32:10 on 28/09/2004. Each sensor samples at 20 Hz, yielding $T=2000$ observations per series.
@@
According to the Northern California Earthquake Catalogue [^1] https://ncedc.org/ncedc/catalog-search.html (NCEC), four microearthquakes (events with magnitude $<$ 2.0) were recorded in Parkfield during this time. Prior to our analysis, we apply a 2-16 Hz Butterworth bandpass filter to each series, as is standard procedure in the seismology literature (Caudron et al.,, 2018). We then normalise each series to have zero mean and unit variance. For the likelihood model within each segment we use the AR model described in Equation (9), setting the lag to $L=1$ and use hyperparameters $\alpha_{j}=\beta_{j}=\delta_{j,1}=1$ for all series $j=1,\dots,7$. We then ran two MCMC chains in parallel for 20,000 iterations using the particle Gibbs scheme with 200 particles, discarding the first 2000 iterations of each chain as burn-in.
@@
Figure 2: Network analysis of seismic sensor data: (a) Heatmap displaying estimated inclusion probabilities $Pr(A_{i,j}=1|\boldsymbol{y})$. (b) Locations of the seven seismic sensors (yellow triangles) and the epicentres of the microearthquakes (red circles) labelled according to their order of occurrence. Arrows correspond to edges obtained after thresholding $Pr(A_{i,j}=1|\boldsymbol{y})$ at 50%.
@@
Figure 2 displays the posterior edge inclusion probabilities and sensor/microearthquake locations. We perform link edge analysis by thresholding $Pr(A_{i,j}=1|\boldsymbol{y})$ at 50%. The resulting edges displayed in Figure 2(b) suggest that sensors nearer to Parkfield tend to lead the ones which are further away.
@@
Figure 3 displays the processed data and posterior change-point probabilities for each station. According to the NCEC, the first microearthquake occurs at 22:32:15.6 and its epicentre is 9km from the closest station - FROB, see Figure 2(b). Our NetCP model detects the first change-point at FROB with a high probability ($\approx 0.78$) at 22:32:17.35, a delay of 1.75 seconds. It is known [^2] See for example: www.bgs.ac.uk/discovering-geology/earth-hazards/earthquakes/how-are-earthquakes-detected/ that P-waves travel fastest in the earths crust, at around 5-7 km/s. Thus, the expected time taken for the P-waves to reach FROB is between 1.29 - 1.8 seconds, consistent with the time of the first change-point under our model. Similarly, the furthest station, SMNB, is located 20.8km from the first epicentre and our model detects the first change-point here at 22:32:19.05, a delay of 3.45 seconds. Again, this is consistent with the expected time of arrival for P-waves (between 2.97 - 4.16 seconds). These delays agree with the inferred posterior graph structure displayed in Figure 2, which indicates that change-points propagate away from the epicentres.
@@
Figure 3: Processed signals (top) and estimated posterior change-point probabilities (bottom) under the proposed model for the seven geophone recording stations. Blue dashed lines indicate the recorded start times for the four microearthquakes.
4.3 EEG Data
@@
We now apply our model to EEG signals collected from paediatric epilepsy patients at Boston Children’s Hospital. An EEG is a medical test which measures the electrical activity of the brain using electrodes placed on the scalp, and is a key tool for diagnosing and monitoring epilepsy. These recordings provide important information about the onset, origin, and spread of seizures across brain regions. Change-point analysis may be used to uncover such information. For example, change-points may first appear in the signals from the temporal lobe electrodes before propagating to signals from frontal electrodes shortly after.
@@
The data was obtained from the CHB-MIT Scalp EEG Database (Guttag,, 2010) and was originally analysed by Shoeb, (1981). We consider EEG data from patient CHB01, recorded during session 21. The raw signals are represented as bipolar montages, calculated as the voltage difference between pairs of adjacent electrodes in order to reduce noise. Temporal Lobe Epilepsy, characterized by seizures originating in the temporal lobes, is one of the most common forms of focal epilepsy in children. Based on this, we limit our analysis to six bipolar montages derived from eight electrodes placed in the temporal and frontal regions, see Figure 4. We focus on a 15 second segment of data from 325 to 340 seconds, during which a seizure is clinically annotated to begin at 327 seconds.
@@
Before we proceed to analysing the EEG data, we preprocess it by applying a 1-30 Hz Butterworth bandpass filter in order to isolate the relevant frequency components (Chung et al.,, 2024). The signals, originally sampled at 256 Hz, are then downsampled by a factor of 3. We then apply first-order differencing and normalise each series to have zero mean and unit variance. The processed dataset then consists of $d=6$ time series, each containing $T=1279$ observations. We then use the same AR likelihood model and hyperparameter settings as in Section 4.2 and again ran two MCMC chains for 20,000 iterations with 200 particles and a burn-in of 2000.
@@
Figure 4: EEG Electrode Positions: (a) Labelled brain lobes coloured by region. (b) Top down view of electrode locations with colours corresponding to those in (a). Blue circles represent electrode locations. Blue edge with red dots indicate the bipolar montages that make up the raw signals. Black arrows indicate the edges obtained after thresholding $Pr(A_{i,j}=1|\boldsymbol{y})$ at 50%.
Figure 5 displays the processed EEG signals and posterior change-point probabilities. The first change-point appears in the T8-P8 signal, and is consistent with the clinically annotated seizure start time. Subsequent change-points can be seen in all or subsets of the signals as the seizure activity continues to evolve and propagate to other regions of the brain.
@@
Figure 4(a) displays the four brain lobes while Figure 4(b) displays the locations of the electrodes and bipolar montages which make up the six signals. Figure 4(b) displays the edges obtained after thresholding the posterior edge inclusion probabilities $Pr(A_{i,j}=1|\boldsymbol{y})$ at 50%. The edges suggest that change-points are first observed in the right temporal lobes (electrodes T8/P8) before spreading to the left temporal lobe (electrodes T7/P7) and to the frontal lobes (electrodes FP1, FP2, F7, F8) shortly after. This suggests that the seizure originated in the region of the P8 electrode and is consistent with previous analysis by Chung et al., (2024), where the EEG signals were visually analysed by two neurologists. Although this dataset does not include diagnostic information about the patient, the observed propagation pathway in Figure 4(a) aligns with patterns commonly associated with Temporal Lobe Epilepsy (see Jacobs et al.,, 2008). Our analysis also supports the concept of epileptogenic networks, where seizures develop and propagate through localised brain networks rather than appearing as isolated events (see Bartolomei et al.,, 2025).
@@
Figure 5: Processed EEG signals (top) and estimated posterior change-point probabilities (bottom) under the proposed model for the six bipolar montages. Blue dashed lines indicate the clinically annotated seizure onset time.
@@ 5 Discussion
Empirical results show that the proposed method outperforms existing approaches in scenarios with strong inter-series dependence in change-point structure. Moreover, the model successfully recovers interpretable latent network structures in real-world seismological and EEG datasets, providing a principled and automated way to detect and quantify lead-lag behaviour previously observed in these domains.
@@
The proposed model may also be extended in several ways. A central assumption is that observations within a segment are independent of those in other segments. However, in many practical settings, this assumption may be overly restrictive—for example, when certain data characteristics persist or evolve gradually across segment boundaries. An extension to incorporate Markov switching dynamics, where segment-specific parameters can transition between a set of recurring states, may provide a more flexible framework for modelling data exhibiting distinct and recurring regimes. Similarly, the latent directed graph is assumed to be fixed. Future work may relax this to allow the graph to change at certain times, reflecting a shift in the lead-lag relationships.
Acknowledgements
Carson McKee was supported by the Heilbronn Institute for Mathematical Research. Data for this study come from the High Resolution Seismic Network (HRSN) doi:10.7932/HRSN, operated by the UC Berkeley Seismological Laboratory, which is archived at the Northern California Earthquake Data Center (NCEDC), doi: 10.7932/NCEDC. The simulation study was aided by the King’s Computational Research, Engineering and Technology Environment (CREATE) computing cluster.
Appendix A: Detailed MCMC Steps

### v2

@@ 4.1 Simulation Study
Table 2: Average model log Bayes Factors relative to NetCP, with 5% quantiles computed over 50 simulated datasets for each change-point scenario under the Gaussian means likelihood. The highest average Bayes Factor in each scenario is shown in bold. Values above zero indicate models outperforming NetCP.
Table 2 displays the averaged log Bayes factors, relative to NetCP, for each model under the Gaussian means likelihood scenario. In S1 and S2, which involve lead-lag relationships in change-points, NetCP clearly outperforms all other methods, with substantial negative log Bayes factors observed for competing models. In S3, where change-points occur simultaneously across series, NG-CCP achieves the highest score relative to NetCP, though the interval overlaps zero, suggesting comparable performance. In S4, G-CCP performs best, though differences between methods are again modest. For S5, which contains independent change-points across series, G-CCP yields the highest average Bayes factor, but Yao performs similarly, as indicated by their overlapping intervals. Overall, these results suggest that NetCP excels in structured, temporally misaligned settings, while G-CCP and Yao are more competitive in scenarios with simultaneous or independent change-points.
@@
Table 3: Average model log Bayes Factors relative to NetCP, with 5% quantiles computed over 50 simulated datasets for each change-point scenario under the AR Process likelihood. The highest average Bayes Factor in each scenario is shown in bold. Values above zero indicate models outperforming NetCP.
@@ 4.2 Seismology Data
We first consider the application of our model to seismology data. We apply the NetCP model to the detection of microearthquakes using data from a network of seismic sensors. Specifically, we use recordings from the High Resolution Seismic Network (HSRN), maintained by the UC Berkeley Seismology Laboratory [80]. This network comprises geophone sensors positioned around the town of Parkfield, California, and is designed to monitor seismic activity along the San Andreas fault. Similar data was previously analyzed by [91] who took a frequentist approach, focusing on estimating the relative delay between change-points in different series. For our analysis, we consider data from seven sensors over a 100 second period beginning at 22:32:10 on 28/09/2004. Each sensor samples at 20 Hz, yielding $T=2000$ observations per series.
@@
According to the Northern California Earthquake Catalog [^1] https://ncedc.org/ncedc/catalog-search.html (NCEC), four microearthquakes (events with magnitude << 2.0) were recorded in Parkfield during this time. Prior to our analysis, we apply a 2-16 Hz Butterworth bandpass filter to each series, as is standard procedure in the seismology literature [52]. We then normalize each series to have zero mean and unit variance. For the likelihood model within each segment we use the AR model described in Equation (9), setting the lag to $L=1$ and use hyperparameters $\alpha_{j}=\beta_{j}=\delta_{j,1}=1$ for all series $j=1,\dots,7$. We then ran two MCMC chains in parallel for 20,000 iterations using the particle Gibbs scheme with 200 particles, discarding the first 2000 iterations of each chain as burn-in.
@@
Figure 4: Network analysis of seismic sensor data: (a) Heatmap displaying estimated inclusion probabilities $Pr(A_{i,j}=1|\boldsymbol{y})$. (b) Locations of the seven seismic sensors (yellow triangles) and the epicenters of the microearthquakes (red circles) labeled according to their order of occurrence. Arrows correspond to edges obtained after thresholding $Pr(A_{i,j}=1|\boldsymbol{y})$ at 50%.
@@
Figure 4 displays the posterior edge inclusion probabilities between sensors alongside the locations of the seven seismic sensors and the microearthquake epicenters. Arrows indicate edges inferred by NetCP after thresholding the inclusion probabilities at 50%. The resulting edges suggest that sensors nearer to Parkfield tend to lead the ones which are further away, with the sensor FROB likely to experience change-points first and SMNB likely to experience them last.
@@
Figure 5 displays the processed data (top) and posterior change-point probabilities (bottom) for each seismic sensor. According to the NCEC, the first microearthquake occurs at 22:32:15.6 and its epicenter is 9km from the closest sensor - FROB, see Figure 4(b). Our NetCP model detects the first change-point at FROB with a high probability ($\approx 0.78$) at 22:32:17.35, a delay of 1.75 seconds. It is known [^2] See for example: www.bgs.ac.uk/discovering-geology/earth-hazards/earthquakes/how-are-earthquakes-detected/ that P-waves travel fastest in the earths crust, at around 5-7 km/s. Thus, the expected time taken for the P-waves to reach FROB is between 1.29 - 1.8 seconds, consistent with the time of the first change-point under our model. Similarly, the furthest sensor, SMNB, is located 20.8km from the first epicenter and our model detects the first change-point here at 22:32:19.05, a delay of 3.45 seconds. Again, this is consistent with the expected time of arrival for P-waves (between 2.97 - 4.16 seconds). These delays agree with the inferred posterior graph structure displayed in Figure 4, which indicates that change-points propagate away from the epicenters.
@@
Figure 5: Processed signals (top) and estimated posterior change-point probabilities in red (bottom) under the proposed model for the seven geophone recording stations. Blue dashed lines indicate the recorded start times for the four microearthquakes.
4.3 Electroencephalogram (EEG) Data
@@
We now apply our model to EEG signals collected from pediatric epilepsy patients at Boston Children’s Hospital. An EEG is a medical test which measures the electrical activity of the brain using electrodes placed on the scalp, and is a key tool for diagnosing and monitoring epilepsy. These recordings provide important information about the onset, origin, and spread of seizures across brain regions. Change-point analysis may be used to uncover such information. For example, change-points may first appear in the signals from the temporal lobe electrodes before propagating to signals from frontal electrodes shortly after.
@@
The data was obtained from the CHB-MIT Scalp EEG Database [68] and was originally analyzed by [86]. We consider EEG data from patient CHB01, recorded during session 21. The raw signals are represented as bipolar montages, calculated as the voltage difference between pairs of adjacent electrodes in order to reduce noise. Temporal Lobe Epilepsy, characterized by seizures originating in the temporal lobes, is one of the most common forms of focal epilepsy in children. Based on this, we limit our analysis to six bipolar montages derived from eight electrodes placed in the temporal and frontal regions, see Figure 6. We focus on a 15 second segment of data from 325 to 340 seconds, during which a seizure is clinically annotated to begin at 327 seconds.
@@
Before we proceed to analyzing the EEG data, we preprocess it by applying a 1-30 Hz Butterworth bandpass filter in order to isolate the relevant frequency components [56]. The signals, originally sampled at 256 Hz, are then downsampled by a factor of 3. We then apply first-order differencing and normalize each series to have zero mean and unit variance. The processed dataset then consists of $d=6$ time series, each containing $T=1279$ observations. We then use the same AR likelihood model and hyperparameter settings as in Section 4.2 and again ran two MCMC chains for 20,000 iterations with 200 particles and a burn-in of 2000.
@@
Figure 6: EEG Electrode Positions: (a) Labeled brain lobes colored by region. (b) Top down view of electrode locations with colors corresponding to those in (a). Blue circles represent electrode locations. Blue edge with red dots indicate the bipolar montages that make up the raw signals. Black arrows indicate the edges obtained after thresholding $Pr(A_{i,j}=1|\boldsymbol{y})$ at 50%.
@@
Figure 6(a) displays the four brain lobes while Figure 6(b) displays the locations of the electrodes and bipolar montages which make up the six signals. Arrows indicate the edges obtained after thresholding the posterior edge inclusion probabilities $Pr(A_{i,j}=1|\boldsymbol{y})$ at 50%. The edges suggest that change-points are first observed in the right temporal lobes (electrodes T8/P8) before spreading to the left temporal lobe (electrodes T7/P7) and to the frontal lobes (electrodes FP1, FP2, F7, F8) shortly after. This suggests that the seizure originated in the region of the P8 electrode and is consistent with previous analysis by [56], where the EEG signals were visually analyzed by two neurologists. Although this dataset does not include diagnostic information about the patient, the observed propagation pathway in Figure 6(a) aligns with patterns commonly associated with Temporal Lobe Epilepsy [[, see]]Jacobs2008. Our analysis also supports the concept of epileptogenic networks, where seizures develop and propagate through localized brain networks rather than appearing as isolated events [[, see]]Bartolomei2025.
Figure 7 displays the processed EEG signals (top) and posterior change-point probabilities (bottom). The first change-point appears in the T8-P8 signal, and is consistent with the clinically annotated seizure start time. Subsequent change-points can be seen in all or subsets of the signals as the seizure activity continues to evolve and propagate to other regions of the brain.
@@
Figure 7: Processed EEG signals (top) and estimated posterior change-point probabilities, in red, (bottom) under the proposed model for the six bipolar montages. Blue dashed lines indicate the clinically annotated seizure onset time.
@@ 5 Discussion
Empirical results show that the proposed method outperforms existing approaches in scenarios with strong inter-series dependence in change-point structure. Moreover, the model successfully recovers interpretable latent network structures in real-world seismological and EEG datasets, providing a principled and automated way to detect and quantify lead-lag behavior previously observed in these domains.
@@
The proposed model may also be extended in several ways. A central assumption is that observations within a segment are independent of those in other segments. However, in some practical settings, this assumption can be restrictive, for example, when certain data characteristics persist or evolve gradually across segment boundaries. An extension to incorporate Markov switching dynamics, where segment-specific parameters can transition between a set of recurring states, could provide a more flexible framework for modeling data that exhibit distinct and recurring regimes. Finally, in this paper we have assumed a fixed latent directed graph. For our future work we plan to relax this assumption and allow the graph to change at certain times, reflecting a shift in the lead-lag relationships.
Acknowledgments
Carson McKee was supported by the Heilbronn Institute for Mathematical Research. The simulation study was aided by the King’s Computational Research, Engineering and Technology Environment (CREATE) computing cluster.
Appendix A: Proofs
Proof of Proposition 1
According to Equation (16), for any time $t>1$, we have $W_{0,j}q_{0,j}/Z_{j}\leq p_{j}(\boldsymbol{x}_{t-1})<1$ and $0<(1-p_{j}(\boldsymbol{x}_{t-1}))\leq 1-W_{0,j}q_{0,j}/Z_{j}$. Hence the chain is irreducible and positive recurrent since there is always positive probability of the chain resetting the run length to 1 or incrementing. The chain is also clearly aperiodic. Convergence to stationarity then follows using standard results from ergodic theory, e.g. Theorem 13.0.1 of [79].
Proof of Proposition 2:
Let $t>1$. First note that since $A_{1,1}=A_{2,1}=0$, then we have $Pr(X_{1,t}=1|\boldsymbol{X}_{t-1}=\boldsymbol{x}_{t-1})=q_{0,1}$ and $Pr(X_{1,t}=x_{1,t-1}+1|\boldsymbol{X}_{t-1}=\boldsymbol{x}_{t-1})=1-q_{0,1}$. It is then easy to verify that
$Pr(X_{1,t}=k)=\begin{cases}q_{0,1}(1-q_{0,1})^{k-1},\;\;\;\text{for }k=1,...,t-1\\ (1-q_{0,1})^{k-1},\;\;\;\;\;\;\;\;\;\text{for }k=t;\end{cases}$
Now, for $t>1$ and $Z_{2}=W_{0,2}+W_{1,2}$:
$Pr(U_{2,t}=1)$ $=$ $\sum_{k=1}^{t-1}Pr(X_{2,t}=1|X_{1,t-1}=k)Pr(X_{1,t-1}=k)$
$=$ $\sum_{k=1}^{t-1}Z_{2}^{-1}\left[W_{0,2}q_{0,2}+W_{1,2}g_{1,2}(k)\boldsymbol{1}(k<t-1)\right]Pr(X_{1,t-1}=k)$
$=$ $Z_{2}^{-1}\left[W_{0,2}q_{0,2}+W_{1,2}\sum_{k=1}^{t-1}g_{1,2}(k)\boldsymbol{1}(k<t-1)Pr(X_{1,t-1}=k)\right]$
$=$ $Z_{2}^{-1}\left[W_{0,2}q_{0,2}+W_{1,2}\lambda(t-1,q_{0,1},q_{1,2})\right]$
where
$\lambda(t,q_{0,1},q_{1,2})$ $=$ $\sum_{k=1}^{t}g_{1,2}(k)\boldsymbol{1}(k<t)Pr(X_{1,t}=k)$
$=$ $\sum_{k=1}^{t-1}g_{1,2}(k)Pr(X_{1,t}=k)$
$=$ $\sum_{k=1}^{t-1}q_{1,2}(1-q_{1,2})^{k-1}q_{0,1}(1-q_{0,1})^{k-1}$
$=$ $q_{0,1}q_{1,2}\sum_{k=0}^{t-2}[(1-q_{0,1})(1-q_{1,2})]^{k}$
$=$ $q_{0,1}q_{1,2}\left[\frac{1-[(1-q_{0,1})(1-q_{1,2})]^{t-1}}{1-(1-q_{0,1})(1-q_{1,2})}\right]$
as required.
Proof of Proposition 3:
Let $t>1$ and $h\geq 1$. We compute
$\mathbb{E}(U_{1,t}U_{2,t+h})$ $=$ $Pr(X_{1,t}=1,X_{2,t+h}=1)$
$=$ $Pr(X_{1,t}=1)Pr(X_{2,t+h}=1|X_{1,t}=1)$
$=$ $q_{0,1}\sum_{k=1}^{t+h-1}Pr(X_{2,t+h}=1|X_{1,t+h-1}=k,X_{1,t}=1)Pr(X_{1,t+h-1}=k|X_{1,t}=1)$
$=$ $q_{0,1}\sum_{k=1}^{t+h-1}Pr(X_{2,t+h}=1|X_{1,t+h-1}=k)Pr(X_{1,t+h-1}=k|X_{1,t}=1)$
Since the change-points in series one follow an independent Bernoulli process, we have $Pr(X_{1,t+h-1}=k|X_{1,t}=1)=Pr(X_{1,h}=k)$ for $k\in\{1,...,h\}$. We also have that $\boldsymbol{1}(X_{1,t+h-1}<t+h-1)=1$. Therefore,
$\mathbb{E}(U_{1,t}U_{2,t+h})$ $=$ $q_{0,1}\sum_{k=1}^{h}Z_{2}^{-1}[W_{0,2}q_{0,2}+W_{1,2}g_{1,2}(k)]Pr(X_{1,h}=k)$
$=$ $q_{0,1}Z_{2}^{-1}[W_{0,2}q_{0,2}+W_{1,2}\sum_{k=1}^{h}g_{1,2}(k)Pr(X_{1,h}=k)]$
Now substituting $g_{1,2}(k)=q_{1,2}(1-q_{1,2})^{k-1}$ and following similar steps as in Proposition 2, we obtain
$\mathbb{E}(U_{1,t}U_{2,t+h})=q_{0,1}Z_{2}^{-1}[W_{0,2}q_{0,2}+W_{1,2}\lambda^{*}(h,q_{0,1},q_{1,2})]$
where
$\lambda^{*}(t,q_{0,1},q_{1,2})=\lambda(t,q_{0,1},q_{1,2})+q_{12}[(1-q_{0,1})(1-q_{1,2})]^{t-1}$
and $\lambda(t,q_{0,1},q_{1,2})$ is given above in Proposition 2. The result then follows combining this with $\mathbb{E}(U_{1,t})=q_{0,1}$ and $\mathbb{E}(U_{2,t+h})=Pr(U_{2,t+h}=1)$ given in Proposition 2.
Appendix B: Detailed MCMC Steps